# 3DGStream: On-the-fly Training of 3D Gaussians for Efficient Streaming of Photo-Realistic Free-Viewpoint Videos

### v1

@@ 1 Introduction
Recently, Kerbl et al. [26] have achieved real-time radiance field rendering using 3D Gaussians (3DGs), thus enabling the instant synthesis of novel views in static scenes with just minutes of training. Inspired by this breakthrough, we propose 3DGStream, a method that utilizes 3DGs to construct Free-Viewpoint Videos (FVVs) of dynamic scenes. Specifically, we first train the initial 3DGs on the multi-view frames at timestep 0. Then, for each timestep $i$, we use the 3DGs of previous timestep $i-1$ as initialization and pass it to a two-stage pipeline. (1) In Stage 1, we train a Neural Transformation Cache (NTC) to model the transformations of 3DGs. (2) Then in the Stage 2, we use an adaptive 3DG addition strategy to handle emerging objects by spawning frame-specific additional 3DGs near these objects and optimize them along with periodic splitting and pruning. After the two-stage pipeline concludes, we use both the 3DGs transformed by the NTC and the additional 3DGs for rendering at the current timestep $i$, with only the former carrying over for initialization of the subsequent timestep. This design significantly reduces the storage requirements for the FVV, as we only need to store the per-frame NTCs and frame-specific additions, rather than all 3DGs for each frame.
@@
  • We conducted extensive experiments to demonstrate 3DGStream’s competitive edge in rendering quality, training time, and requisite storage, as well as its superior rendering speed, compared to existing state-of-the-art dynamic scene reconstruction methods.
Figure 3: Overview of 3DGStream. Given a set of multi-view video streams, 3DGStream aims to construct high-quality FVV stream of the captured dynamic scene on-the-fly. Initially, we optimize a set of 3DGs to represent the scene at timestep 00. For each subsequent timestep $i$, we use the 3DGs from timestep $i-1$ as an initialization and then engage in a two-stage training process: Stage 1: We train the Neural Transformation Cache (NTC) to model the translations and rotations of 3DGs. After training, the NTC transforms the 3DGs, preparing them for the next timestep and the next stage in the current timestep. Stage 2: We spawn frame-specific additional 3DGs at potential locations and optimize them along with periodic splitting and pruning. After the two-stage process concludes, both transformed and additional 3DGs are used to render at the current timestep i, with only the transformed ones carried into the next timestep.
@@ 2.1 Novel View Synthesis for Static Scenes
Synthesizing novel views from a set of images of static scenes is a time-honored problem in the domains of computer vision and graphics. Traditional methods such as Lumigraph [22, 8] or Light-Field [28, 50, 10, 16] achieve new view synthesis through interpolation. In recent years, Neural Radiance Fields (NeRF) [36] have achieved photorealistic synthesizing results by representing the radiance field using a multi-layer perceptron (MLP). A series of subsequent works enhance NeRF’s performance in various aspects, such as accelerating training speeds [52, 40, 12, 20, 25, 13], achieving real-time rendering [64, 21, 14, 72, 23, 47], and improving synthesis quality on challenging scenes [35, 2, 3, 56, 37, 4] or sparse inputs [73, 11, 41, 69, 63, 66, 53]. Since the vanilla NeRF employs costly volume rendering, necessitating neural network queries for rendering, subsequent approaches faced trade-offs in training time, rendering speed, model storage, image quality, and applicability. To address these challenges, Kerbl et al. [26] propose 3D Gaussian Splatting (3DG-S), which integrates of 3DGs with differentiable point-based rendering. 3DG-S enables real-time high-fidelity view synthesis in large-scale unbounded scenes after brief training periods with modest storage requirements. Inspired by this work, we extend its application to the task of constructing FVVs of dynamic scenes. Taking it a step further, we design a on-the-fly training framework to achieve efficient FVV streaming.
@@ 2.2 Free-Viewpoint Videos of Dynamic Scenes
Prior-driven methods [62, 74, 30, 68, 27] leverage parametric models or incorporate additional priors, such as human-pose skeletons, to bolster performance on the reconstruction of specific dynamic objects, e.g., humans. However, their application is limited and not generalizable to broader scenes.
@@
Warp-based methods [55, 42, 44, 46, 51, 1, 61] assumpt that the dynamics of the scene arise from the deformation of static structures. These techniques warp the radiance field of each frame onto one or several canonical frames, achieving notable results. However, the strong assumptions they rely on often prevent them from handling topological variations.
@@
Among the aforementioned works, only NeRFPlayer [51], ReRF [59], StreamRF [29], and Dynamic3DG [34] are able to stream FVVs. NeRFPlayer achieves FVV streaming through a decomposition module and a feature streaming module, but it is only able to stream pre-trained models. ReRF and Dynamic3DG are limited to processing inward-facing scenes that are equipped with foreground masks and contain only a few objects, necessitating several minutes to train each frame. StreamRF stands out by requiring only a few seconds for each frame’s training to construct high-fidelity FVVs on challenging real-world dynamic scenes with compressed model storage. However, it falls short in rendering speed. Contrarily, our approach matches or surpasses StreamRF in training speed, model storage, and image quality, all while achieving real-time rendering at 200 FPS.
@@ 2.3 Concurrent Works
Except for Dynamic3DG, several concurrent works have extended 3DG-S to represent dynamic scenes. Deformable3DG [70] employs an MLP to model the deformation of 3DGs, while [65] introduces a hexplane-based encoder to enhance the efficency of deformation query. Meanwhile, [71, 18] lift 3DG to 4DG primitives for dynamic scene representation. However, these approaches are limited to offline reconstruction and lack streamable capabilities, whereas our work aims to achieve efficient streaming of FFVs with an online training paradigm.
@@ 4 Method
3DGStream constructs photo-realistic FVV streams from multi-view video streams on-the-fly using a per-frame training paradigm. We initiate the process by training 3DGs [26] at timestep 0. For subsequent timesteps, we employ the previous timestep’s 3DGs as an initialization and pass them to a two-stage pipeline. Firstly (Sec. 4.1), a Neural Transformation Cache (NTC) is trained to model the transformation for each 3DG. Once the training is finished, we transform the 3DGs and carry the transformed 3DGs to the next timestep. Secondly (Sec. 4.2), we employ an adaptive 3DG addition strategy to handle emerging objects. For each FVV frame, we render views at the current timestep using both the transformed 3DGs and additional 3DGs, while the latter are not passed to the next timestep. Note that we only need to train and store the parameters of the NTC and the additional 3DGs for each subsequent timestep, not all the 3DGs. We depict an overview of our approach in Fig. 3.
4.1 Neural Transformation Cache
For NTC, we seek a structure that is compact, efficient, and adaptive to model the transformations of 3DGs. Compactness is essential to reduce the model storage. Efficiency enhances training and inference speeds. Adaptivity ensures the model focuses more on dynamic regions. Additionally, it would be beneficial if the structure could consider certain priors of dynamic scenes [24, 5, 54], such as the tendency for neighboring parts of an object to have similar motion.
Inspired by Neural Radiance Caching [39] and I-NGP [40], we employ multi-resolution hash encoding combined with a shallow fully-fused MLP [38] as the NTC. Specifically, following I-NGP, we use multi-resolution voxel grids to represent the scene. Voxel grids at each resolution are mapped to a hash table storing a $d$-dimensional learnable feature vector. For a given 3D position $x\in\mathbb{R}^{3}$, its hash encoding at resolution $l$, denoted as $h(x;l)\in\mathbb{R}^{d}$, is the linear interpolation of the feature vectors corresponding to the eight corners of the surrounding grid. Consequently, its multi-resolution hash encoding $h(x)=[h(x;0),h(x;1),...,h(x;L-1)]\in\mathbb{R}^{Ld}$, where $L$ represents the number of resolution levels. The multi-resolution hash encoding addresses all our requirements for the NTC:
  • Compactness: Hashing effectively reduces the storage space needed for encoding the whole scene.
  • Efficiency: Hash table lookup operates in $O(1)$, and is highly compatible with modern GPUs.
  • Adaptivity: Hash collisions occur in hash tables at finer resolutions, allowing regions with larger gradients—representing dynamic regions in our context—to drive the optimization.
  • Priors: The combination of linear interpolation and the voxel-grid structure ensures the local smoothness of transformations. Additionally, the multi-resolution approach adeptly merges global and local information.
Furthermore, to enhance the performance of NTC with minimal overhead, we utilize a highly-optimized shallow fully-fused MLP [38]. This maps the hash encoding to a 7-dimensional output: the first three dimensions indicate the translation of the 3DG; the remaining dimensions represent the rotation of the 3DG using quaternions. Given multi-resolution hash encoding coupled with MLP, our NTC is formalized as:
${d\mu,dq}=MLP(h(\mu)),$ (5)
where $\mu$ denotes the mean of the input 3DG. We transform the 3DGs based on $d\mu$ and $dq$. Specifically, the following parameters of the transformed 3DGs are given as:
  • Mean: $\mu^{\prime}=\mu+d\mu$, where $\mu^{\prime}$ is the new mean and $+$ represents vector addition.
  • Rotation: $q^{\prime}=norm(q)\times norm(dq)$, where $q^{\prime}$ is the new roation, $\times$ indicates quaternion multiplication and $norm$ denotes normalization.
  • SH Coefficients: Upon rotating the 3DG, the SH coefficients should also be adjusted to align with the rotation of the 3DG. Leveraging the rotation invariance of SH, we directly employ SH Rotation to update SHs. Please refer to the supplementary materials (Suppl.) for details.
In Stage 1, we transform the 3DGs from the previous frame by NTC and then render with them. The parameters of the NTC is optimized by the loss between the rendered image and the ground truth. Following 3DG-S [26], the loss function is $L_{1}$ combined with a D-SSIM term:
$L=(1-\lambda)L_{1}+\lambda L_{D-SSIM},$ (6)
where $\lambda=0.2$ in all our experiments. It should be noted that during the training process, the 3DGs from the previous frame remain frozen and do not undergo any updates. This implies that the input to the NTC remains consistent.
Additionally, to ensure training stability, we initialize the NTC with warm-up parameters. The loss employed during the warm-up is defined as:
$L_{warm-up}=||d\mu||_{1}-cos^{2}(norm(dq),Q),$ (7)
where $Q$ is the identity quaternion. The first term uses the $L_{1}$ norm to ensure the estimated translation approaches zero, while the second term, leveraging cosine similarity, ensures the estimated rotation approximates no rotation. However, given the double-covering property of the unit quaternions, we use the square of the cosine similarity. For each scene, we execute the warm-up solely after the training at timestep 0, using noise-augmented means of the initial 3DGs as input. After 3000 iterations of training (roughly 20 seconds), the parameters are stored and used to initialize the NTCs for all the following timesteps.
(a) I-NGP [40]
(b) HyperReel [1]
(c) StreamRF [29]
(d) 3DGStream
(e) Ground Truth
Figure 4: Qualitative comparisons on the discussion scene of the Meet Room dataset and the sear steak scene of the N3DV dataset.
4.2 Adaptive 3DG Addition
Relying solely on 3DGs transformations adequately cover a significant portion of real-world dynamic scenes, with translations effectively managing occlusions and disappearances in subsequent timesteps. However, this approach falters when faced with objects not present in the initial frame, such as transient objects like flames or smoke, and new persistent objects like the liquid poured out of a bottle. Since 3DG is an unstructured explicit representation, it’s essential to add new 3DGs to model these emerging objects. Considering constraints related to model storage requirements and training complexities, it’s not feasible to generate an extensive number of additional 3DGs nor allow them to be used in subsequent frames, as this would cause 3DGs to accumulate over time. This necessitates a strategy for swiftly generating a limited number of frame-specific 3DGs to model these emerging objects precisely and thereby enhance the completeness of the scene at the current timestep.
Firstly, we need to ascertain the locations for the emerging objects. Inspired by 3DG-S [26], we recognized the view-space positional gradients of 3DGs as a key indicator. We observed that for emerging objects, the 3DGs in proximity exhibited large view-space positional gradients. This is attributed to the optimization attempting to ‘masquerade’ the emerging object by transforming the 3DGs. However, since we prevent the colors of the 3DGs from being updated in Stage 1, this attempt falls short. Nonetheless, they are still transformed to appropriate positions, with large view-space positional gradients.
Based on the aforementioned observations, we deem it appropriate to introduce additional 3DGs around these high-gradient regions. Moreover, to exhaustively capture every potential location where new objects might emerge, we adopt an adaptive 3DG spawn strategy. Specifically, we track view-space positional gradient during the final training epoch of Stage 1. Once this stage concludes, we select 3DGs that have an average magnitude of view-space position gradients exceeding a relatively low threshold $\tau_{grad}=0.00015$. For each selected 3DG, the position of the additional 3DG is sampled from $X\sim\mathcal{N}(\mu,2\Sigma)$, where $\mu$ and $\Sigma$ is the mean and the convariance matrix of the selected 3DG. While we avoid assumptions about the other attributes of the additional 3DGs, improper initializations of SH coefficients and scaling vectors tend to result in an optimization preference for reducing opacity over adjusting these parameters. This causes additional 3DGs to quickly become transparent, thereby failing to capture the emerging objects. To mitigate this issue, the SH coefficients and scaling vectors of these 3DGs are derived from the selected ones, with rotations set to the identity quaternion q = [1, 0, 0, 0] and opacity initialized at 0.1. After spawning, the 3DGs undergo optimization utilizing the same loss function (Eq. 6) as Stage 1. Note that only the parameters of the additional 3DGs are optimized, while those of the transformed 3DGs remain fixed.
To guard against local minima and manage the number of additional 3DGs, we implement an adaptive 3DG quantity control strategy. Specifically, in Stage 2, we set a relatively high threshold, $\tau_{\alpha}=0.01$, for the opacity value. At the end of each training epoch, for 3DGs with view-space position gradients exceeding $\tau_{grad}$, we spawn additional 3DGs nearby to address under-reconstructed regions. These additional 3DGs inherit their rotations and SH coefficients from the original 3DG, but their scaling is adjusted to 80% of the original, mirroring the ‘split’ operation described by Kerbl et al. [26]. Subsequently, we discard any additional 3DGs with opacity values below $\tau_{\alpha}$ to suppress the growth in the quantity of 3DGs.
@@ 4.2 Adaptive 3DG Addition
Offline DyNeRF [31] ${}^{\dagger}$ 0.1 260 0.02 $\times$
@@
Online StreamRF [29] 30.68 ${}^{\star}$ 0.25 8.3 $\checkmark$
Ours 31.67 ${}^{\star}$ 0.20 215 $\checkmark$
Table 1: Quantitative comparison on the N3DV dataset. The training time, required storage and PSNR are averaged over the whole 300 frames for each scene. ${}^{\dagger}$yNeRF [31] only report metrics on the flame salmon scene. ${}^{\star}$onsidering the initial model.
@@
StreamRF [29] 26.72 ${}^{\star}$ 0.17 10
Ours 30.79 ${}^{\star}$ 0.10 288
Table 2: Quantitative comparison on the Meet Room dataset. Note that the training time, required storage and PSNR are averaged over the whole 300 frames. ${}^{\star}$onsidering the initial model.
@@ 5.3 Comparisons
Qualitative comparisons. While our approach primarily aims to enhance the efficiency of online FVV construction, as illustrated in Tabs. 1 and 2, it still achieves competitive image quality. In Fig. 4, we present a qualitative comparison with I-NGP [40], HyperReel [1], and StreamRF [29] across scenes on the N3DV dataset [31] and the Meet Room dataset [29], with a special emphasis on dynamic objects such as faces, hands, and tongs, as well as intricate objects like labels and statues. It is evident that our method faithfully captures the dynamics of the scene without sacrificing the ability to reconstruct intricate objects. Please refer to our project page for more video results.
@@ 7 Conclusion
We propose 3DGStream, an novel method for efficient Free-Viewpoint Video streaming. Based on 3DG-S [26], we utilizes an effective Neural Transformation Cache to capture the motion of objects. In addition, we propose an Adaptive 3DG Addition strategy to accurately model emerging objects in dynamic scenes. The two-stage pipeline of 3DGStream enables the on-the-fly reconstruction of dynamic scenes in video streams. While ensuring photo-realistic image quality, 3DGStream achieves on-the-fly training ($\sim$10s per-frame) and real-time rendering ($\sim$200FPS) at megapixel resolution with moderate requisite storage. Extensive experiments demonstrate the efficiency and effectiveness of 3DGStream.
8 Acknowledgement
This work was supported in part by Zhejiang Province Program (2022C01222, 2023C03199, 2023C03201), the National Program of China (62172365, 2021YFF0900604, 19ZDA197), Ningbo Science and Technology Plan Project (022Z167, 2023Z137), and MOE Frontier Science Center for Brain Science & Brain-Machine Integration (Zhejiang University).

### v3

@@ 1 Introduction
Recently, Kerbl et al. [24] have achieved real-time radiance field rendering using 3D Gaussians (3DGs), thus enabling the instant synthesis of novel views in static scenes with just minutes of training. Inspired by this breakthrough, we propose 3DGStream, a method that utilizes 3DGs to construct Free-Viewpoint Videos (FVVs) of dynamic scenes. Due to procedural reasons, we are currently unable to show the text of our method on arXiv. If you are interested in the implementation, please refer to the source code on our project page or check out the paper published on CVPR Open Access.
@@
  • We conduct extensive experiments to demonstrate 3DGStream’s competitive edge in rendering quality, training time, and requisite storage, as well as its superior rendering speed, compared to existing state-of-the-art dynamic scene reconstruction methods.
@@ 2.1 Novel View Synthesis for Static Scenes
Synthesizing novel views from a set of images of static scenes is a time-honored problem in the domains of computer vision and graphics. Traditional methods such as Lumigraph [21, 7] or Light-Field [26, 47, 9, 15] achieve new view synthesis through interpolation. In recent years, Neural Radiance Fields (NeRF) [34] has achieved photorealistic synthesizing results by representing the radiance field using a multi-layer perceptron (MLP). A series of subsequent works enhance NeRF’s performance in various aspects, such as accelerating training speeds [49, 37, 11, 19, 23, 12], achieving real-time rendering [60, 20, 13, 68, 22, 44], and improving synthesis quality on challenging scenes [33, 2, 3, 52, 35, 4] or sparse inputs [69, 10, 38, 65, 59, 62, 50]. Since the vanilla NeRF employs costly volume rendering, necessitating neural network queries for rendering, subsequent approaches faced trade-offs in training time, rendering speed, model storage, image quality, and applicability. To address these challenges, Kerbl et al. [24] propose 3D Gaussian Splatting (3DG-S), which integrates of 3DGs with differentiable point-based rendering. 3DG-S enables real-time high-fidelity view synthesis in large-scale unbounded scenes after brief training periods with modest storage requirements. Inspired by this work, we extend its application to the task of constructing FVVs of dynamic scenes. Taking it a step further, we design a on-the-fly training framework to achieve efficient FVV streaming.
@@ 2.2 Free-Viewpoint Videos of Dynamic Scenes
Prior-driven methods [58, 70, 28, 64, 25] leverage parametric models or incorporate additional priors, such as skeletons, to bolster performance on the reconstruction of specific dynamic objects, e.g., humans. However, their application is limited and not generalizable to broader scenes.
@@
Warp-based methods [51, 39, 41, 43, 48, 1, 57] assumpt that the dynamics of the scene arise from the deformation of static structures. These methods warp the radiance field of each frame onto one or several canonical frames, achieving notable results. However, the strong assumptions they rely on often prevent them from handling topological variations.
@@
Among the aforementioned works, only NeRFPlayer [48], ReRF [55], StreamRF [27], and Dynamic3DG [32] are able to stream FVVs. NeRFPlayer achieves FVV streaming through a decomposition module and a feature streaming module, but it is only able to stream pre-trained models. ReRF and Dynamic3DG are limited to processing scenes with few objects and foreground mask, necessitating minute-level per-frame training times. StreamRF stands out by requiring only a few seconds for each frame’s training to construct high-fidelity FVVs on challenging real-world dynamic scenes with compressed model storage. However, it falls short in rendering speed. Contrarily, our approach matches or surpasses StreamRF in training speed, model storage, and image quality, all while achieving real-time rendering at 200 FPS.
@@ 2.3 Concurrent Works
Except for Dynamic3DG, several concurrent works have extended 3DG-S to represent dynamic scenes. Deformable3DG [66] employs an MLP to model the deformation of 3DGs, while [61] introduces a hexplane-based encoder to enhance the efficency of deformation query. Meanwhile, [67, 17] lift 3DG to 4DG primitives for dynamic scene representation. However, these approaches are limited to offline reconstruction and lack streamable capabilities, whereas our work aims to achieve efficient streaming of FVVs with an online training paradigm.
@@ 4 Method
Due to procedural reasons, we are currently unable to show the text of our method on arXiv. If you are interested in the implementation, please refer to the source code on our project page or check out the paper published on CVPR Open Access.
@@
Offline DyNeRF [29] 29.58 0.1 260 0.02 $\times$
@@
Online StreamRF [27] 30.68 17.7/31.4 0.25 8.3 $\checkmark$
Ours 31.67 7.6/7.8 0.20 215 $\checkmark$
Table 1: Quantitative comparison on the N3DV dataset. The training time, required storage and PSNR are averaged over the whole 300 frames for each scene. DyNeRF [29] only report metrics on the flame salmon scene. Considering the initial model.
@@
StreamRF [27] 26.72 5.7/9.0 0.17 10
Ours 30.79 4.0/4.1 0.10 288
Table 2: Quantitative comparison on the Meet Room dataset. Note that the training time, required storage and PSNR are averaged over the whole 300 frames. Considering the initial model.
@@ 5.3 Comparisons
Qualitative comparisons. While our approach primarily aims to enhance the efficiency of online FVV construction, as illustrated in Tabs. 1 and 2, it still achieves competitive image quality. In LABEL:fig:visual_Comparisons, we present a qualitative comparison with I-NGP [37], HyperReel [1], and StreamRF [27] across scenes on the N3DV dataset [29] and the Meet Room dataset [27], with a special emphasis on dynamic objects such as faces, hands, and tongs, as well as intricate objects like labels and statues. It is evident that our method faithfully captures the dynamics of the scene without sacrificing the ability to reconstruct intricate objects. Please refer to our project page for more video results.
@@ 7 Conclusion
We propose 3DGStream, an novel method for efficient Free-Viewpoint Video streaming. Based on 3DG-S [24], we utilizes an effective Neural Transformation Cache to capture the motion of objects. In addition, we propose an Adaptive 3DG Addition strategy to accurately model emerging objects in dynamic scenes. The two-stage pipeline of 3DGStream enables the online reconstruction of dynamic scenes in video streams. While ensuring photo-realistic image quality, 3DGStream achieves on-the-fly training ($\sim$10s per-frame) and real-time rendering ($\sim$200FPS) at megapixel resolution with moderate requisite storage.